OpenStudy (anonymous):

Which of the following is the same as 2 × 2 × 2 × 2 × 2 × 2? Select one of the options below as your answer: A. 2^6 B. 2^7 C. 6^2 D. 64^6

OpenStudy (anonymous):

There are 6 of them. So, it is 2^6.

OpenStudy (anonymous):

Raising a number to a power means multiplying that number with itself as many times as the power it is raised to. a^n = a x a x a x a x ....... x a (n times).

OpenStudy (apoorvk):

so it means 2 is multiplied 6 times with itself. And any no. multiplied with itself n times is the no. raised to the power of n \[(a*a*a*a*a...*a)_{n times} = a^n\]

OpenStudy (apoorvk):

So what do you think the answer would be? Is GT right? ;)